If x is 11 1/9 % percent more than y, then y is what percent less than x?
A 9% b 10% c 11 1/9% d 12 1/2%

Answer by dabanfield(803)

If x is 11 1/9 % percent more than y, then y is what percent less than x?
A 9% b 10% c 11 1/9% d 12 1/2%
Note that 11 1/9 % = 1/9. So we have:
x = y + (1/9)*y
x = (9/9)*y + (1/9)*y
x = (10/9)*y
y = (9/10)*x
y then is 90% of x so y is 10% less than x.